OpenStudy (anonymous):

if F(x)=1/x+2 what is F(3+h) ????

OpenStudy (anonymous):

i have no clue. i dont even know where to start

OpenStudy (calculusfunctions):

Is it\[f(x)=\frac{ 1 }{ x }+2\]OR\[f(x) =\frac{ 1 }{ x +2 }\]Either way f(3 + h) requires you substitute 3 + h for x in f(x).

OpenStudy (anonymous):

the second one

OpenStudy (calculusfunctions):

Alright so then go ahead and perform the substitution like I said.

OpenStudy (anonymous):

i honestly dont know what i am doing :(

OpenStudy (michele_laino):

@kimmykaht please, perform this substitution: x=h+3, in your function f(x)

OpenStudy (calculusfunctions):

Do you understand? For example, suppose we have some function\[f(x) =\frac{ x }{ x +3 }\]And we require f(1 + a). Then\[f(1+a)=\frac{ 1+a }{ 1+a +3 }\]\[f(1+a)=\frac{ 1+a }{ 4+a }\]You see all I did there was replace every x in the original equation with 1 + a. Now you try!!

OpenStudy (anonymous):

am i solving for h?

OpenStudy (anonymous):

h=1-15f?

OpenStudy (calculusfunctions):

NO! Did I solve for a in my example??

OpenStudy (anonymous):

... geez sorry

OpenStudy (anonymous):

im obviously not intelligent

OpenStudy (anonymous):

math is hard for me, that is why i am seeking help

OpenStudy (calculusfunctions):

No need to apologize. I wasn't angry or implying a lack of patience. I don't think that you're not intelligent.

OpenStudy (anonymous):

so what did i do wrong? if im not suppose to solve for h, what do i do?

OpenStudy (calculusfunctions):

Did you look at my example and try to follow it?

OpenStudy (anonymous):

never mind, i figured it out... its \[\frac{ 1 }{ h+5 }\]

OpenStudy (calculusfunctions):

Yes! That is exactly what I have been trying to convey to you.